OpenStudy (anonymous):

john owns a hotdog stand His Profit is represented by P(x)=-x^2+20x-10 with P(x) being profit and x the number of hot dogs sold. What is the most he can earn?

OpenStudy (tkhunny):

Does the expression "-b/(2a)" look familiar?

OpenStudy (anonymous):

yes it does

OpenStudy (tkhunny):

This will tell you the value of 'x' where P(x) is at a maximum. x = -20/(2(-1)) = -20/(-2) = 10 Find P(10) and you will have the maximum. This answers the question.